A rectangular water tank is open at the top. Its capacity is 24m3. Its length and breadth are 4m and 3m respectively. Ignoring the thickness of the material used for the building the tank, the total cost of painting the inner and outer surfaces of the tank at Rs. 10 per m2 ,is
A. Rs. 400
B. Rs. 500
C. Rs. 600
D. Rs. 800

Grade:12th pass

2 Answers

Pawan Prajapati
2 years ago
Our expert is working on this Class X Maths answer. We will update the answer very soon.
Partha
one year ago
L=4m,b=3m,let h =xm
Capacity of tank is 24 m^3
i.e volume=l*b*h=24
=>4*3*x= 24
=>12x=24
X=2m=h.
Since the tank is open & thickness of material is to be ignored.
So we have sum of inner & outer surface 
=2(2lh+2bh+lb)
=2(2*4*2+2*3*2+4*3)
=2(16+12+12)
=80 m^3.
Cost of painting per m^2=rs10
Total cost = 80*10= 800.ans is D-800.